# On the norms of the multiplication operators between weighted Bergman spaces

### 17 Mar 2026 v1

@@ 1. Introduction and main results
Let $\Delta=\{z:|z|<1\}$ denote the unit disk in the complex plane $\mathbb{C}$. For an open domain $\Omega$ in $\mathbb{C}$, we use $\mathcal{A}(\Omega)$ to denote the class of all analytic functions in $\Omega$. For $1<\gamma<\infty$, we define the weighted Bergman space ${\bf A}_{\gamma}^{2}={\bf A}_{\gamma}^{2}(\Delta)$ as
${\bf A}_{\gamma}^{2}(\Delta)=\{\phi\in\mathcal{A}(\Delta):\|\phi\|_{\gamma}^{2}:=(\gamma+1)\iint_{\Delta}|\phi(z)|^{2}(1-|z|^{2})^{\gamma}\frac{dxdy}{\pi}<\infty\}.$
@@
For a locally univalent function $f$ in an open domain $\Omega$ of $\mathbb{C}$, the Schwarzian derivative $S(f)$ of $f$ which is defined by
$S_{f}(z)=\bigg[\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\bigg]^{\prime}-\frac{1}{2}\left[\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\right]^{2},\,z\in\Omega.$
@@
(1.1) $\beta_{f}(t):=\limsup_{r\rightarrow 1^{-}}\frac{\log\int_{-\pi}^{\pi}|f^{\prime}(re^{i\theta})|^{t}|d\theta}{|\log(1-r)|}.$
@@ 2. Some Auxiliary Results
${\bf{A}}_{\gamma}^{2}(\mathbb{C}_{+})=\left\{h\in\mathcal{A}(\mathbb{C}_{+}):\|h\|_{\alpha}^{2}=\iint_{\mathbb{C}_{+}}|h(u+iv)|^{2}v^{\alpha}\frac{dudv}{\pi}<\infty\right\}.$
@@
Similarly, from 2.8, we have
@@
Choose $\delta=\min\{r_{0},r_{2}\}$. Note that for any $\tau$ with $|\tau|<\delta$ and for any $z\in\mathbb{D}$ and $t\in[0,1]$, the points $a_{0}z$ and $a_{0}z+t\tau z$ lie in $\overline{\Delta}_{r_{0}}$ because
@@
we see that, for any given $\varepsilon>0$, there is a constant $r_{1}\in(0,1)$ such that
(2.16) ${\bf Int_{1}}:=\frac{\beta+1}{\pi}\int_{|z|>r_{0}}|g(az)\phi(z)-g(e^{i\theta}z)\phi(z)|^{2}(1-|z|^{2})^{\beta}dxdy<\varepsilon.$
@@
From (2.16) and (2.18), we conclude that
@@
for any $\theta\in[0,2\pi)$. This proves Sub-Lemma 2.7.
Finally, for any $\theta\in\mathbb{R}$, it is easy to check that $\|\phi(e^{i\theta}z)\|_{\alpha}=\|\phi(z)\|_{\alpha}$ for any $\phi\in\mathbf{A}_{\alpha}^{2}$. Then we can obtain by a change of variables that
@@
for any $\theta\in[0,2\pi)$. On the other hand, by means of Sub-Lemma 2.6 and 2.7, we know that, for a fixed $\phi\in\mathbf{A}_{\alpha}^{2}$, the mapping
@@
In view of this fact, we shall consider the following three three subclasses of $\mathcal{R}$.
@@
Here, $P_{n}$ and $P_{m}$ have no common factors except $1$, and $P_{m}(z)$ has no zeros in $\overline{\Delta}=\Delta\cup\mathbb{T}$. It is easy to see that $\mathcal{U}_{P}\subset\mathcal{L}_{I}.$
@@
Now, to prove Theorem 1.5, it is enough to show that following result holds.
@@
Assume that $R$ is a locally univalent function in $\Delta$ with having at least one critical point on $\mathbb{T}$. We first show that $S(R)$ is a multiplier from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\alpha+4}^{2}$.
When $R$ belongs to $\mathcal{Q}_{I}$, as in 3.8, we have
@@
(3.11) $S(R)(z)=-\frac{3}{2}\sum_{r=1}^{s}\frac{1}{(z-z_{r})^{2}}+\sum_{i=1}^{s}\frac{A_{r}(z)}{z-z_{s}}+{\bf Anal}(z).$
@@
When $R$ belongs to $\mathcal{Q}_{II}$, as in 3.9, we have
@@
Here, ${\bf Anal}(z)$ is defined as in (3.12). for each $r\in[1,s]$, $j\in[1,l]$, $B_{i}(z)$ and $C_{j}(z)$ are analytic in $\overline{\Delta}$.
When $R$ belongs to $\mathcal{Q}_{III}$, as in 3.10, we have
@@
Here, ${\bf Anal}(z)$ is the same as in (3.12) and $D_{r},E_{j},F_{k}$ are all analytic in $\Delta$ for each $r\in[1,s],j\in[1,l],k\in[1,t]$. Note that when $t=0$, there is no the following terms
$2\sum_{k=1}^{t}\frac{1}{(z-e^{i\widetilde{\theta}_{k}})^{2}}+\sum_{k=1}^{t}\frac{F_{k}(z)}{z-e^{i\theta_{j}}},$
in the right side of 3. Combining (3.11), (3) and (3.11), we can see that $S(R)$ is a multiplier from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\alpha+4}^{2}$.
@@
here, ${\pi}_{1}$ is the sector $\{z=re^{i\theta}:r\in[0,1],\theta\in[-\eta_{1}-\Theta,\eta_{2}+\Theta]$. Since 1z1-z is continuous in the compact set ${\bf\pi}_{1}$ and the point $1$ does not belongs to ${\bf\pi}_{1}$, then we see that there is a constant $C_{1}>0$ such that
@@
here, ${\bf\pi}_{2}$ is the sector $\{z=re^{i\theta}:r\in[0,1],\theta\in[-\eta_{2},\eta_{2}]$. As $e^{-i\Theta}-z$ is continuous in the compact set $\pi_{2}$ and the point $e^{-i\Theta}$ is not contained in $\pi_{2}$, hence we know that there is a constant $C_{2}>0$ such that
@@
[…]
here, $\pi_{3}$ is the following compact set $\pi_{3}=\overline{\Delta}-s_{1}-s_{2}$, where
@@
$\|S(R)(re^{i\arg z_{1}}z)(1-rz)^{-\lambda}\|_{\alpha+4}^{2}\geq\|-\frac{3}{2}\frac{z_{1}^{-2}{\bf\Pi}(z)}{(1-rz)^{2}}\|_{\alpha+4}^{2}-2\bf\Sigma.$
@@
(4.2) $\|g_{0}\phi\|_{\beta}^{2}\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+2)}{\Gamma(\alpha+2)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}\|\phi\|_{\alpha}^{2}.$
@@
$\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+1)}{\Gamma(\alpha+1)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}\iint_{\Delta}|\phi(z)|^{2}(1-|z|^{2})^{\alpha}dxdy.$
@@
(4.3) $\iint_{\mathbb{C}_{+}}\frac{|h(w)|^{2}}{|w|^{\beta-\alpha}}v^{\beta}dudv\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+1)}{\Gamma(\alpha+1)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}\iint_{\mathbb{C}_{+}}|h(w)|^{2}v^{\alpha}dudv.$
@@
$\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+1)}{\Gamma(\alpha+1)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}\frac{\Gamma(\alpha+1)}{2^{\alpha}}\int_{0}^{\infty}|\tilde{h}(t)|^{2}\frac{dt}{t^{\alpha+1}}.$
@@
$\leq\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(\beta/2+1)}\Big]^{2}\int_{0}^{\infty}|\tilde{h}(t)|^{2}\frac{dt}{t^{\alpha+1}}.$

### 18 Mar 2026 v2

@@ 1. Introduction and main results
Let $\Delta=\{z:|z|<1\}$ denote the unit disk in the complex plane $\mathbb{C}$. For an open domain $\Omega$ in $\mathbb{C}$, we use $\mathcal{A}(\Omega)$ to denote the class of all analytic functions in $\Omega$. For $1<\alpha<\infty$, we define the weighted Bergman space ${\bf A}_{\alpha}^{2}={\bf A}_{\alpha}^{2}(\Delta)$ as
${\bf A}_{\alpha}^{2}(\Delta)=\{\phi\in\mathcal{A}(\Delta):\|\phi\|_{\alpha}^{2}:=(\alpha+1)\iint_{\Delta}|\phi(z)|^{2}(1-|z|^{2})^{\alpha}\frac{dxdy}{\pi}<\infty\}.$
@@
For a locally univalent function $f$ in an open domain $\Omega$ of $\mathbb{C}$, the Schwarzian derivative $S(f)$ of $f$ is defined by
$S(f)(z)=\bigg[\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\bigg]^{\prime}-\frac{1}{2}\left[\frac{f^{\prime\prime}(z)}{f^{\prime}(z)}\right]^{2},\,z\in\Omega.$
@@
(1.1) $\beta_{f}(t):=\limsup_{r\rightarrow 1^{-}}\frac{\log\int_{-\pi}^{\pi}|f^{\prime}(re^{i\theta})|^{t}d\theta}{|\log(1-r)|}.$
@@ 2. Some Auxiliary Results
${\bf{A}}_{\alpha}^{2}(\mathbb{C}_{+})=\left\{h\in\mathcal{A}(\mathbb{C}_{+}):\|h\|_{\alpha}^{2}=\iint_{\mathbb{C}_{+}}|h(u+iv)|^{2}v^{\alpha}\frac{dudv}{\pi}<\infty\right\}.$
@@
Similarly, from (2.8), we have
@@
Choose $\delta=\min\{r_{0},r_{2}\}$. Note that for any $\tau$ with $|\tau|<\delta$ and for any $z\in\mathbb{D}$ and $t\in[0,1]$, the points $a_{0}z$ and $a_{0}z+t\tau z$ lie in $\overline{\Delta}_{r_{1}}$ because
@@
we see that, for any given $\varepsilon>0$, there is a constant $r_{0}\in(0,1)$ such that
${\bf Int_{1}}:=\frac{\beta+1}{\pi}\int_{|z|>r_{0}}|g(az)\phi(z)-g(e^{i\theta}z)\phi(z)|^{2}(1-|z|^{2})^{\beta}dxdy<\varepsilon.$
@@
From (2) and (2.17), we conclude that
@@
for any $\theta\in(-\pi,\pi]$. This proves Sub-Lemma 2.7.
Finally, for any $\theta\in(-\pi,\pi]$, it is easy to check that $\|\phi(e^{i\theta}z)\|_{\alpha}=\|\phi(z)\|_{\alpha}$ for any $\phi\in\mathbf{A}_{\alpha}^{2}$. Then we can obtain by a change of variables that
@@
for any $\theta\in(-\pi,\pi]$. On the other hand, by means of Sub-Lemma 2.6 and 2.7, we know that, for a fixed $\phi\in\mathbf{A}_{\alpha}^{2}$, the mapping
@@
In view of this fact, we shall consider the following three subclasses of $\mathcal{R}$.
@@
Here, $P_{n}$ and $P_{m}$ have no common factors except $1$, and $P_{m}(z)$ has no zeros in $\overline{\Delta}=\Delta\cup\mathbb{T}$.
@@
Now, to prove Theorem 1.5, it is enough to show that the following result holds.
@@
Assume that $R$ is a locally univalent function in $\Delta$ having at least one critical point on $\mathbb{T}$. We first show that $S(R)$ is a multiplier from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\alpha+4}^{2}$.
When $R$ belongs to $\mathcal{Q}_{I}$, as in (3.8), we have
@@
(3.11) $S(R)(z)=-\frac{3}{2}\sum_{r=1}^{s}\frac{1}{(z-z_{r})^{2}}+\sum_{r=1}^{s}\frac{A_{r}(z)}{z-z_{r}}+{\bf Anal}(z).$
@@
When $R$ belongs to $\mathcal{Q}_{II}$, as in (3.9), we have
@@
Here, ${\bf Anal}(z)$ is defined as in (3.12). for each $r\in[1,s]$, $j\in[1,l]$, $B_{r}(z)$ and $C_{j}(z)$ are analytic in $\overline{\Delta}$.
When $R$ belongs to $\mathcal{Q}_{III}$, as in (3.10), we have
@@
Here, ${\bf Anal}(z)$ is the same as in (3.12) and $D_{r},E_{j},F_{k}$ are all analytic in $\Delta$ for each $r\in[1,s],j\in[1,l],k\in[1,t]$. Note that when $t=0$, there are no following terms
$2\sum_{k=1}^{t}\frac{1}{(z-e^{i\widetilde{\theta}_{k}})^{2}}+\sum_{k=1}^{t}\frac{F_{k}(z)}{z-e^{i\widetilde{\theta}_{k}}},$
in the right side of (3). Combining (3.11), (3) and (3), we can see that $S(R)$ is a multiplier from ${\bf A}_{\alpha}^{2}$ to ${\bf A}_{\alpha+4}^{2}$.
@@
here, ${\pi}_{1}$ is the sector $\{z=re^{i\theta}:r\in[0,1],\theta\in[-\eta_{1}-\Theta,\eta_{1}+\Theta]\}$. Since 1z1-z is continuous in the compact set ${\bf\pi}_{1}$ and the point $1$ does not belong to ${\bf\pi}_{1}$, then we see that there is a constant $C_{1}>0$ such that
@@
here, ${\bf\pi}_{2}$ is the sector $\{z=re^{i\theta}:r\in[0,1],\theta\in[-\eta_{2},\eta_{2}]\}$. As $e^{-i\Theta}-z$ is continuous in the compact set $\pi_{2}$ and the point $e^{-i\Theta}$ is not contained in $\pi_{2}$, hence we know that there is a constant $C_{2}>0$ such that
@@
$\sup_{r\in(0,1),z\in\Delta-\Omega_{1}-\Omega_{2}}|1-rz|^{-2\lambda-2}|1-re^{i\Theta}z|^{-2}=\max_{z\in\pi_{3}}\{|1-z|^{-2\lambda-2}|e^{-i\Theta}-z|^{-2}\}$
here, $\pi_{3}$ is the following compact set $\pi_{3}=\overline{\Delta}-(s_{1}\cup s_{2})$, where
@@
$\|S(R)(re^{i\arg z_{1}}z)(1-rz)^{-\lambda}\|_{\alpha+4}^{2}\geq\|-\frac{3}{2}\frac{z_{1}^{-2}{\bf\Pi}(z)}{(1-rz)^{2}}\|_{\alpha+4}^{2}-4\bf\Sigma.$
@@
(4.2) $\|g_{0}\phi\|_{\beta}^{2}\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+2)}{\Gamma(\alpha+2)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(1+\beta/2)}\Big]^{2}\|\phi\|_{\alpha}^{2}.$
@@
$\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+1)}{\Gamma(\alpha+1)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(1+\beta/2)}\Big]^{2}\iint_{\Delta}|\phi(z)|^{2}(1-|z|^{2})^{\alpha}dxdy.$
@@
(4.3) $\iint_{\mathbb{C}_{+}}\frac{|h(w)|^{2}}{|w|^{\beta-\alpha}}v^{\beta}dudv\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+1)}{\Gamma(\alpha+1)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(1+\beta/2)}\Big]^{2}\iint_{\mathbb{C}_{+}}|h(w)|^{2}v^{\alpha}dudv.$
@@
$\leq\frac{1}{2^{\beta-\alpha}}\frac{\Gamma(\beta+1)}{\Gamma(\alpha+1)}\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(1+\beta/2)}\Big]^{2}\frac{\Gamma(\alpha+1)}{2^{\alpha}}\int_{0}^{\infty}|\tilde{h}(t)|^{2}\frac{dt}{t^{\alpha+1}}.$
@@
$\leq\Big[\frac{\Gamma(1+\alpha/2)}{\Gamma(1+\beta/2)}\Big]^{2}\int_{0}^{\infty}|\tilde{h}(t)|^{2}\frac{dt}{t^{\alpha+1}}.$